OpenStudy (kayders1997):

F'(t)=2cost+sec^2t -pi/2

OpenStudy (kayders1997):

@zepdrix

zepdrix (zepdrix):

\[\large\rm f'(t)=2\cos t+\sec^2t,\qquad\qquad f\left(\frac{\pi}{3}\right)=4\]So what are you stuck on? :o

OpenStudy (kayders1997):

I think I did the f(t) right

OpenStudy (kayders1997):

F(t)=2sint+tan(t)+c?

OpenStudy (kayders1997):

Is that right? :O

zepdrix (zepdrix):

k good

zepdrix (zepdrix):

So we have a "general solution".\[\large\rm f(t)=2\sin t+\tan t+c\]They gave us `initial data` which we can use to figure out the unknown constant. This will give us a "particular solution"

OpenStudy (kayders1997):

:O

OpenStudy (kayders1997):

So does it have to equal 4?

OpenStudy (kayders1997):

Hmmmm no?

OpenStudy (anonymous):

it has to equal 4 when \(t=\frac{\pi}{3}\)

OpenStudy (kayders1997):

So the t will be pi/3 and the whole thing equals 4?

OpenStudy (anonymous):

yes, then you can solve for \(C\)

OpenStudy (kayders1997):

Okay so

OpenStudy (kayders1997):

Square root of 3+1=4

OpenStudy (anonymous):

\[\large\rm f(t)=2\sin t+\tan t+c\\ \large\rm f(\frac{\pi}{3})=2\sin(\frac{\pi}{3})+\tan (\frac{\pi}{3})+c=4\]

OpenStudy (anonymous):

not quite, but close \[\tan(\frac{\pi}{3})\neq 1\]

OpenStudy (kayders1997):

It doesn't equal 1? :O

OpenStudy (kayders1997):

Oh yeah ughhhh

OpenStudy (kayders1997):

Square root of 3

OpenStudy (kayders1997):

So square root 3 + square root of 3 = 2square roots of 3

OpenStudy (kayders1997):

So .539

OpenStudy (kayders1997):

Can you help me with one more?